Calculate the payoff at expiration for a put holder on an option on a Eurodollar future, where the underlying IMM index value at expiration using the IMM quotation on a 90-day dollar denominated time deposit on a $1,000,000 notional principal is 98.64 and the “exercise price” (also as an IMM index value) is 98.80

$0

$250

$300

$350

$400
Answer #1

Put option holder has right to sell the asset at exercise price

Profit on exercise = 98.8-98.64 = 0.16%

=1000000*0.16%*90/360

=$400